OpenStudy (anonymous):

A force of ten pounds compresses a 21 inch spring 8 inches. how much work is done in compressing the spring from a length of 15 inches to a length of 13 inches?

OpenStudy (anonymous):

\[\int\limits_{13}^{15}(\frac{ 5 }{ 4 } x) dx\] ???

OpenStudy (anonymous):

its not though the answer is 17.5 ft-lb... that gives 35 ft-lb, what am i doing wrong

OpenStudy (anonymous):

sorry 35 in-lb not ft-lb

OpenStudy (perl):

We know from hook's law F = k*d 10 = k*8 10/8 = k 5/4 = k

OpenStudy (anonymous):

yeah hookes law used it got that 5/4=k, I am stuck I have no clue how to get to 17.5 ft lb from that

OpenStudy (perl):

you have to consider the natural length of the spring

OpenStudy (anonymous):

so do I subtract an integral for 15 to 21?

OpenStudy (perl):

$$ \Large \int_{21-15}^{21-13}\frac{ 5 }{ 4 } x~ dx $$

OpenStudy (anonymous):

lol awesome I was pulling my hair out, that will not be forgotten... So is that 17.5 now in ft lb or still in inch lb?

OpenStudy (perl):

we are in inch*lbs because of how we found k (using inches and pounds)

OpenStudy (anonymous):

thanks you perl

OpenStudy (perl):

$$ \Large \int_{21-15}^{21-13}\frac{ 5 }{ 4 } x~ dx = \Large \int_{6}^{8}\frac{ 5 }{ 4 } x~ dx = 17.5 ~~inch \cdot lbs $$